# Cyber security of OT networks: A tutorial and overview

### v1

@@ 5 Regulatory Frameworks & Compliance in OT Cybersecurity
As the cybersecurity landscape for OT networks evolves, regulatory frameworks and compliance measures play a crucial role in the protection of critical infrastructure, industrial automation, and national security [49]. Given the increasing sophistication of cyber threats, the implementation of standardized security guidelines has become most important in mitigating risks to industrial control systems (ICS) such as SCADA, PLCs, and DCS [50]. These frameworks ensure a well-structured risk assessment, strong security policies, and a concrete approach to OT resilience.
With the growing integration of IT-OT convergence and the rising risks associated with cyber-physical systems, organizations are facing increased challenges in navigating a complex regulatory landscape [51]. This section examines the most influential global cybersecurity standards, recent developments in compliance enforcement, and innovative approaches to strengthening regulatory security.
@@ 5.1 NIST Cybersecurity Framework (CSF) & NIST 800-82
The National Institute of Standards and Technology (NIST) has been at the forefront of cybersecurity research and framework development, providing industries with a structured methodology for securing critical infrastructure and ICS environments. Through initiatives like the NIST Cybersecurity Framework (CSF) and its guidance on ICS security, NIST has established itself as a global leader in promoting best practices. These frameworks not only help organizations identify and manage risks, but also support them in aligning their security strategies with evolving threats and compliance requirements. By continuously updating its standards, NIST ensures that industries are equipped to tackle emerging challenges in the ever-changing cybersecurity landscape [52] [53].
@@
    NIST CSF – A Risk-Based Approach to Industrial Security: In the past decade, the NIST CSF has emerged as a foundation of industrial security, offering a structured and adaptable approach to managing cybersecurity risks. NIST CSF was initially released in 2014 and has undergone significant revisions, culminating in the 2024 update [54]. These revisions incorporate modern security paradigms such as:
@@
      Supply Chain Risk Management (SCRM): Supply chain attacks such as the SolarWinds (late 2020) breach [55] and the Kaseya ransomware attack (2021) [56] have demonstrated vulnerabilities in third-party integrations. The 2024 NIST update mandates continuous vendor evaluation, zero-trust mechanisms for software dependencies, and dynamic risk profiling to mitigate risks introduced by suppliers. Organizations must implement hardware attestation to verify firmware integrity and conduct real-time AI-powered anomaly detection in supplier transactions.
@@
      Zero-Trust Architectures (ZTA): Traditional perimeter-based defenses, which rely on securing a defined network boundary, are no longer sufficient in today’s highly interconnected and cloud-driven environments [57]. Cyber threats often bypass these defenses, exploiting risks within trusted networks. To address this, the Zero-Trust model shifts the focus from perimeter security to a framework where trust is never assumed, regardless of whether a user or device is inside or outside the network. The NIST update emphasizes several critical components of ZTA:
@@
        Identity-Based Access Controls (IBAC): Access is granted dynamically based on verified user and machine identities. This ensures that only authenticated and authorized entities can access specific resources, reducing the risk of unauthorized access [58].
@@
        Multi-Layer Authentication Protocols: Advanced methods like FIDO2 (Fast Identity Online) and certificate-based authentication add an extra layer of security. These methods are resistant to phishing attacks and ensure that credentials cannot be easily compromised [59].
@@
        Policy Enforcement Points (PEPs): PEPs act as gatekeepers that validate and enforce access policies in real-time. They continuously monitor and verify every request, ensuring that users and devices comply with predefined security rules before granting access to resources [60].
@@
      ZTA also prioritize continuous monitoring and risk-based adaptive controls, meaning that security is not a one-time check but an ongoing process [36] [61]. This approach strengthens the ability to detect and prevent unauthorized activities, even in complex hybrid IT-OT environments, making it an essential strategy for modern cybersecurity frameworks. This emphasis on eliminating implicit trust and enforcing strict identity verification has driven the widespread adoption of Zero Trust strategies, which have grown significantly from 24% in 2021 to 61% in 2023 [62] [63]. This significant growth shows the adoption of ZTA’s effectiveness in addressing modern security challenges and adapting to evolving cyber threats.
@@
    NIST SP 800-82 – Industrial Control Systems Security: NIST Special Publication 800-82 provides a comprehensive framework for securing ICS in sectors such as manufacturing, energy, water utilities, and critical infrastructure. As ICS environments increasingly integrate with IT networks, the security risks have expanded, necessitating updated defense strategies. The latest edition of NIST SP 800-82 (Revision 3, 2023) emphasizes enhanced security controls that address modern cyber threats, attack vectors, and mitigation techniques [64].
@@
      Network Segmentation and Isolation: Network segmentation is a fundamental principle in ICS cybersecurity, reducing lateral movement of threats between IT and OT environments. NIST 800-82 outlines best practices for implementing segmentation using [65]:
@@
        Industrial Demilitarized Zones (IDMZs): Separating enterprise IT from OT networks while allowing controlled communication [66].
@@
        Unidirectional Security Gateways: Hardware-enforced one-way data flows that prevent external threats from accessing OT systems [67].
@@
        Air-Gapped Systems: Isolating critical OT assets from all external networks to prevent cyber infiltration [68].
@@
      Behavioral Anomaly Detection and AI-driven Threat Intelligence: Traditional signature-based Intrusion Detection Systems (IDS) have proven insufficient against zero-day vulnerabilities and ICS-specific malware. The updated NIST framework encourages ML driven anomaly detection to recognize deviations from normal system behaviors. The AI driven threat monitoring detects real-time deviations in PLCs and SCADA systems. Some real-world incidents like Triton (2017) [69] and Industroyer (2016) [70] demonstrated that ICS malware can manipulate control logic without triggering traditional alarms [71].
@@ 5.2 IEC 62443 – Global Standard for Industrial Automation Security
    IEC 62443-2-1:2024Security Program Requirements for IACS Asset Owners: This updated standard specifies policy and procedure requirements for asset owners to establish and maintain an effective security program for industrial automation systems. The 2024 edition introduces significant technical changes, including a revised requirement structure into Security Program Elements (SPEs), elimination of duplications with Information Security Management Systems (ISMS), and the definition of a maturity model for evaluating requirements [72].
@@
    IEC 62443-3-3System Security Requirements and Security Levels: This part outlines system-level security requirements and defines various security levels to guide asset owners and integrators in implementing appropriate security measures. It emphasizes foundational requirements such as identification and authentication control, use control, system integrity, data confidentiality, and more [73].
@@
    IEC 62443-4-2Technical Security Requirements for IACS Components: This section provides detailed technical security requirements for individual IACS components, ensuring that each component meets specific security capabilities. It aligns with the foundational requirements detailed in IEC 62443-3-3 and emphasizes the importance of secure product development practices [73].
@@
IEC 62443 for Next-Generation Industrial Security: The increasing convergence of OT and IT in modern industrial systems has necessitated the adaptation of IEC 62443 to evolving cybersecurity challenges. The rapid increase of cloud-integrated SCADA systems, AI-driven intrusion detection, and Industrial Internet of Things (IIoT) frameworks has expanded the attack surface of critical infrastructures, demanding a more adaptive and resilient cybersecurity framework. While IEC 62443 remains the de facto global standard for Industrial Automation and Control Systems (IACS) security, its application must account for emerging threats posed by interconnected industrial environments [74]. The integration of ZTA, behavioral anomaly detection, and Secure Access Service Edge (SASE) models into IEC 62443 compliance strategies has become increasingly relevant in mitigating advanced persistent threats (APTs). Notably, ZTNA frameworks ensure granular access control for remote Programmable Logic Controller (PLC) and Remote Terminal Unit (RTU) configurations, minimizing the risks associated with unauthorized network infiltration [75]. Additionally, AI-based Security Information and Event Management (SIEM) platforms are facilitating automated compliance validation, thereby reducing human intervention and improving continuous monitoring efficiency.

### v2

@@ 6 Regulatory Frameworks & Compliance in OT Cybersecurity
As the cybersecurity landscape for OT networks evolves, regulatory frameworks and compliance measures play a crucial role in the protection of critical infrastructure, industrial automation, and national security [bellamkonda2020cybersecurity]. Given the increasing sophistication of cyber threats, the implementation of standardized security guidelines has become most important in mitigating risks to industrial control systems (ICS) such as SCADA, PLCs, and DCS [fortunato2020risk]. These frameworks ensure a well-structured risk assessment, strong security policies, and a concrete approach to OT resilience.
With the growing integration of IT-OT convergence and the rising risks associated with cyber-physical systems, organizations are facing increased challenges in navigating a complex regulatory landscape [qiu2020edge]. This section examines the most influential global cybersecurity standards, recent developments in compliance enforcement, and innovative approaches to strengthening regulatory security.
@@ 6.1 NIST Cybersecurity Framework (CSF) & NIST 800-82
The National Institute of Standards and Technology (NIST) has been at the forefront of cybersecurity research and framework development, providing industries with a structured methodology for securing critical infrastructure and ICS environments. Through initiatives like the NIST Cybersecurity Framework (CSF) and its guidance on ICS security, NIST has established itself as a global leader in promoting best practices. These frameworks not only help organizations identify and manage risks, but also support them in aligning their security strategies with evolving threats and compliance requirements. By continuously updating its standards, NIST ensures that industries are equipped to tackle emerging challenges in the ever-changing cybersecurity landscape [lanz2024updated] [hatinen2024evolution].
@@
    NIST CSF – A Risk-Based Approach to Industrial Security: In the past decade, the NIST CSF has emerged as a foundation of industrial security, offering a structured and adaptable approach to managing cybersecurity risks. NIST CSF was initially released in 2014 and has undergone significant revisions, culminating in the 2024 update [varol2024enhancing]. These revisions incorporate modern security paradigms such as:
@@
      Supply Chain Risk Management (SCRM): Supply chain attacks such as the SolarWinds (late 2020) breach [martinez2021software] and the Kaseya ransomware attack (2021) [robinson2022new] have demonstrated vulnerabilities in third-party integrations. The 2024 NIST update mandates continuous vendor evaluation, zero-trust mechanisms for software dependencies, and dynamic risk profiling to mitigate risks introduced by suppliers. Organizations must implement hardware attestation to verify firmware integrity and conduct real-time AI-powered anomaly detection in supplier transactions.
@@
      Zero-Trust Architectures (ZTA): Traditional perimeter-based defenses, which rely on securing a defined network boundary, are no longer sufficient in today’s highly interconnected and cloud-driven environments [damaraju2024implementing]. Cyber threats often bypass these defenses, exploiting risks within trusted networks. To address this, the Zero-Trust model shifts the focus from perimeter security to a framework where trust is never assumed, regardless of whether a user or device is inside or outside the network. The NIST update emphasizes several critical components of ZTA:
@@
        Identity-Based Access Controls (IBAC): Access is granted dynamically based on verified user and machine identities. This ensures that only authenticated and authorized entities can access specific resources, reducing the risk of unauthorized access [almuseelem2024continuous].
@@
        Multi-Layer Authentication Protocols: Advanced methods like FIDO2 (Fast Identity Online) and certificate-based authentication add an extra layer of security. These methods are resistant to phishing attacks and ensure that credentials cannot be easily compromised [xun2025building].
@@
        Policy Enforcement Points (PEPs): PEPs act as gatekeepers that validate and enforce access policies in real-time. They continuously monitor and verify every request, ensuring that users and devices comply with predefined security rules before granting access to resources [fernandez2024critical].
@@
      ZTA also prioritize continuous monitoring and risk-based adaptive controls, meaning that security is not a one-time check but an ongoing process [syed2022zero] [sample2022zta]. This approach strengthens the ability to detect and prevent unauthorized activities, even in complex hybrid IT-OT environments, making it an essential strategy for modern cybersecurity frameworks. This emphasis on eliminating implicit trust and enforcing strict identity verification has driven the widespread adoption of Zero Trust strategies, which have grown significantly from 24% in 2021 to 61% in 2023 [csoonline] [OKTA]. This significant growth shows the adoption of ZTA’s effectiveness in addressing modern security challenges and adapting to evolving cyber threats.
@@
    NIST SP 800-82 – Industrial Control Systems Security: NIST Special Publication 800-82 provides a comprehensive framework for securing ICS in sectors such as manufacturing, energy, water utilities, and critical infrastructure. As ICS environments increasingly integrate with IT networks, the security risks have expanded, necessitating updated defense strategies. The latest edition of NIST SP 800-82 (Revision 3, 2023) emphasizes enhanced security controls that address modern cyber threats, attack vectors, and mitigation techniques [o2024fiscal].
@@
      Network Segmentation and Isolation: Network segmentation is a fundamental principle in ICS cybersecurity, reducing lateral movement of threats between IT and OT environments. NIST 800-82 outlines best practices for implementing segmentation using [kallatsa2024strategies]:
@@
        Industrial Demilitarized Zones (IDMZs): Separating enterprise IT from OT networks while allowing controlled communication [mazur2016defining].
@@
        Unidirectional Security Gateways: Hardware-enforced one-way data flows that prevent external threats from accessing OT systems [heo2016design].
@@
        Air-Gapped Systems: Isolating critical OT assets from all external networks to prevent cyber infiltration [guri2023air].
@@
      Behavioral Anomaly Detection and AI-driven Threat Intelligence: Traditional signature-based Intrusion Detection Systems (IDS) have proven insufficient against zero-day vulnerabilities and ICS-specific malware. The updated NIST framework encourages ML driven anomaly detection to recognize deviations from normal system behaviors. The AI driven threat monitoring detects real-time deviations in PLCs and SCADA systems. Some real-world incidents like Triton (2017) [myung2019ics] and Industroyer (2016) [makrakis2021vulnerabilities] demonstrated that ICS malware can manipulate control logic without triggering traditional alarms [firoozjaei2022evaluation].
@@ 6.2 IEC 62443 – Global Standard for Industrial Automation Security
    IEC 62443-2-1:2024Security Program Requirements for IACS Asset Owners: This updated standard specifies policy and procedure requirements for asset owners to establish and maintain an effective security program for industrial automation systems. The 2024 edition introduces significant technical changes, including a revised requirement structure into Security Program Elements (SPEs), elimination of duplications with Information Security Management Systems (ISMS), and the definition of a maturity model for evaluating requirements [iec1].
@@
    IEC 62443-3-3System Security Requirements and Security Levels: This part outlines system-level security requirements and defines various security levels to guide asset owners and integrators in implementing appropriate security measures. It emphasizes foundational requirements such as identification and authentication control, use control, system integrity, data confidentiality, and more [ISA].
@@
    IEC 62443-4-2Technical Security Requirements for IACS Components: This section provides detailed technical security requirements for individual IACS components, ensuring that each component meets specific security capabilities. It aligns with the foundational requirements detailed in IEC 62443-3-3 and emphasizes the importance of secure product development practices [ISA].
@@
IEC 62443 for Next-Generation Industrial Security: The increasing convergence of OT and IT in modern industrial systems has necessitated the adaptation of IEC 62443 to evolving cybersecurity challenges. The rapid increase of cloud-integrated SCADA systems, AI-driven intrusion detection, and Industrial Internet of Things (IIoT) frameworks has expanded the attack surface of critical infrastructures, demanding a more adaptive and resilient cybersecurity framework. While IEC 62443 remains the de facto global standard for Industrial Automation and Control Systems (IACS) security, its application must account for emerging threats posed by interconnected industrial environments [leander2019applicability]. The integration of ZTA, behavioral anomaly detection, and Secure Access Service Edge (SASE) models into IEC 62443 compliance strategies has become increasingly relevant in mitigating advanced persistent threats (APTs). Notably, ZTNA frameworks ensure granular access control for remote Programmable Logic Controller (PLC) and Remote Terminal Unit (RTU) configurations, minimizing the risks associated with unauthorized network infiltration [gottel2023qualitative]. Additionally, AI-based Security Information and Event Management (SIEM) platforms are facilitating automated compliance validation, thereby reducing human intervention and improving continuous monitoring efficiency.